I need help find the vertex for the equation 4x+y^2+3y=-5

Answer by lwsshak3(11628)

find the vertex for the equation
4x+y^2+3y=-5
complete the square:
4x+(y^2+3y+9/4)=-5+9/4
4x+(y+3/2)^2=-20/4+9/4=-11/4
4x=-(y+3/2)^2-11/4
x=-(1/4)(y+3/2)^2-11/16
vertex form of equation for parabola that opens leftward: x=-A(y-k)^2+h, (h,k)=(x,y) coordinates of the vertex.
For given problem:
A=-(1/4)
vertex:(-11/16,-3/2)